# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v1

@@ 4 Stabilization of modulus field with SUSY breaking effects
the superpotential for modulus should take the following form
$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}}.$ (22)
for the Kahler potential of the form
@@
in the unit $M_{Pl}=1$ and $c_{1}M_{Pl}$ denotes the typical scale for the superpotential for dimensionless number $c_{1}$.
@@
where ${\cal P}$ denotes a polynomial with respect to $j(\tau)$. The Klein j-function is defined as a modular form of zero weight
@@
We should note that, in such a stabilization mechanism, fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although may not be the global minimum. With proper SUSY breaking contributions, possibly from the ”matter”-like sector, the stabilized modulus value, which located at/near the fixed points $z_{FP}=\omega,i$, can slightly deviate from the original global minimum and possibly uplift the vacuum to dS (de Sitter) one. In fact, the classical 10d supergravity Maldacena-Nunez no-go result MN suggests that any constructions that lead to dS vacuum must include non-classical ingredients as D-branes, O-planes, or quantum corrections. To describe the scenario in which SUSY is linearly realized in a gravitational sector and non-linearly realized in the uplifting SUSY breaking sector, (for example, from anti-D branes) the formalism of non-linearly realized SUSY as formulated by KS can be used for such a system. It was proved in Burgess:2021juk with the non-linear SUSY realization scheme that the supergravity form is stable even if the supermultiplets in SUSY breaking sector are split badly enough to allow some of its members to be integrated out while the others are not.
@@
When SUSY is spontaneously broken in supergravity, the goldstino is eaten by the gravitino via the super-Higgs mechanism. When such a breaking happens at low energies compared with the Planck mass, the goldstino couplings can be described by introducing a (constrained) independent superfield in the supergravity effective action. The goldstino provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. The emergence of the constraints may also be viewed as an effective field theory limit in which certain mass parameters become very large and the corresponding modes decouple.
@@
Coupling of non-linearly realized SUSY with nilpotent goldstino $X_{nl}^{2}=0$ can be described by
${\cal L}\supseteq\int d^{4}\theta X_{nl}^{*}X_{nl}+\left[\int d^{2}\theta f_{0}X_{nl}+h.c.\right]~{},$ (26)
with the chiral superfield
$X_{nl}=\frac{\psi_{X}\psi_{X}}{2F_{X}}+\sqrt{2}\theta\psi_{X}+\theta^{2}F_{X}~{},$ (27)
@@
is the IR limit of certain microscopic superfield. As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $X$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent goldstino. From the perspective of low energy effective theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for modulus stablization and vacuum uplifting discussions.
@@
When coupled to supergravity, the most general form is given by
@@
$K={X}^{*}_{nl}X_{nl}~{},~{}~{}~{}W=f_{0}X+W_{0}~{}.$ (29)
and $f_{0}=-F_{X}$.
To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
$K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{{X}^{*}_{nl}X_{nl}}{(2\Im\tau)^{k_{X}}}~{},$
$W$ $\supseteq$ $\left(f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
The scalar potential for $X$ and $\tau$ can be obtained from
@@
with
@@
$K^{\bar{\alpha}\beta}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{|X|^{2}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}X}{3}\\ -\frac{i(2\Im\tau)k_{X}{X}^{*}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}|X|^{2}}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
Here the indices $\alpha\sim(\tau,X_{nl})$ and the nilpotency of $X_{nl}$ is used in the calculation.
The scalar potential can be calculated to be
@@
and the constraint $X_{nl}^{2}=0$ is used to eliminate some of the high order $X_{nl}$ terms.
We also use the expression for the derivative of Dedekind $\eta$ function, which is given by
@@
The generic complicate form of the scalar potential may obscure the discussions of the stabilization mechanism. So, we would like to concentrate on its typical simplification forms
@@
    The scalar potential can be simplified to be
@@
    -- $\left.3\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\right\}~{}$
    $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}c_{1}^{3}\right|^{2}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{},$
    with
@@
    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{},$ (42)
    Note the constraints $X_{nl}^{2}=\bar{X}_{nl}^{2}=0$ are imposed for the scalar potential.
    The modulus VEV can be determined by the extrema conditions of the scalar potential
@@
    and the minimum condition can be determined by positive definiteness of Hessian matrix in terms of
@@
    with $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$, as well as those conditions for $X_{nl}$. It was argued in target:space ; Ding:2024neh that, modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensures that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
@@
    Then we will determine the extrema condition
@@
    for $X_{nl}\equiv\rho e^{i\theta}$. It can be calculated that the extreme conditions for $X_{nl}$ reduce to the constraints
    $\frac{\partial}{\partial\theta}V=0~{},\Rightarrow\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{},$ (46)
    and
    $\frac{\partial}{\partial\rho}V=0~{},$$\Rightarrow$ $\left|H(\tau)\right|^{2}f_{0}^{2}=0~{},$ (49)
    $\Rightarrow$ $\left\{\begin{array}[]{l}\left|H(\tau)\right|^{2}=0,~{}~{}~{}~{}~{}~{}{\rm for}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}f_{0}\neq 0~{}.\\ ~{}~{}~{}~{}~{}~{}~{}f_{0}=0~{},~{}~{}~{}~{}{\rm and/or}~{}~{}~{}\left|H(\tau)\right|^{2}\neq 0.\end{array}\right.$
    So, we have the minimum conditions for $X_{nl}$
    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{},$ (50)
    in the case $f_{0}\neq 0$. Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes. So, the presence non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe.
    We will concentrate on the constraint $H(\tau)=0$ when $f_{0}\neq 0$. When $f_{0}=0$, the scalar potential for modulus fields will not be affected by $X_{nl}$ as the two sector decouples in the $k_{X}=0$ limit.
    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires either the condition $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0$ or ${\cal P}(j(\tau))=0$. Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively. As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, arbitrary complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all. So, we would like to focus on the constraint
@@
    that satisfies $H(\tau)=0$. Such a constraint reduces to
@@
    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
    So $H^{\prime}(\tau)=0$ can trivially hold for $H(\tau)=0$.
@@
      So, unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$. Similarly, we can deduce that $m\in 2\mathbb{Z}$ should be negative. Such a deduction can also hold for $m\neq 0$ with $n\neq 0$ case, which also requires $m=n=0$. As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
@@
      With the constraints that $m=n=0$, the condition (53) reduces to
@@
      as a constant. As $H(\tau)=0$, the constant should be zero.
    So, the condition (53) requires $H(\tau)$ to be zero. As the constraint (53) is trivial, we would like to discuss the condition $j^{\prime}(\tau)=0$. From the derivative of the Klein j-function
@@
    we can see that its solutions are $\omega$ and $i$.
    So, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential can only be determined by $H(\tau)=0$ (and the extrema conditions from $X_{nl}$ are also satisfied by $z_{0}=\omega,i$). The corresponding vacuum is Minkowski type. As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are the local extrema. Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$. So, other minimum, which is determined by (44) to locate at the boundary of the fundamental domain or at the imaginary axis, is also possible.
    In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices. It can be seen that the energies are non-negative when $f_{0}\neq 0$.
    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$.
@@
    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the . constraint $X_{nl}^{2}=0$. The scalar potential can be simplified to

### v3

@@ 4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
the superpotential for the modulus should take the form
$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}},$ (22)
with the Kahler potential
@@
in the unit $M_{Pl}=1$. Here, $c_{1}M_{Pl}$ denotes the typical mass scale for the superpotential, where $c_{1}$ is a dimensionless number.
@@
where ${\cal P}\left(j(\tau)\right)$ denotes a polynomial with respect to $j(\tau)$. The Klein $j$-function is defined as a modular form of zero weight
@@
We should note that, in such a stabilization mechanism, the fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although they may not be the global minimum. With proper SUSY breaking contributions, possibly from the ”matter”-like sector, the stabilized modulus value, which located at/near the fixed points $z_{FP}=\omega,i$, can slightly deviate from the original global minimum. The minimum of the scalar potential can also possibly be uplifted to dS (de Sitter) type. In fact, the classical 10d supergravity Maldacena-Nunez no-go result MN suggests that any constructions leading to a dS vacuum must include non-classical ingredients, such as D-branes, O-planes or quantum corrections. To describe the scenario in which SUSY is linearly realized in a gravitational sector and non-linearly realized in the uplifting SUSY breaking sector (for example, from anti-D branes), the formalism of non-linearly realized SUSY as formulated by KS can be used for such a system. It was proved in Burgess:2021juk that, with the non-linear SUSY realization scheme, the supergravity form is stable even if the supermultiplets in SUSY breaking sector are split badly enough to allow some of its members to be integrated out while the others are not.
@@
When SUSY is spontaneously broken in supergravity, the Goldstino is eaten by the gravitino via the super-Higgs mechanism. When such a breaking happens at low energies compared with the Planck scale, the Goldstino couplings can be conveniently described by a (constrained) independent superfield in the supergravity effective action. The Goldstino provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. The emergence of the constraints may also be viewed as an effective field theory limit in which certain mass parameters become very large and the corresponding modes decouple.
@@
The coupling of non-linearly realized SUSY with a nilpotent chiral Goldstino superfield, subjected to constraints ${\textbf{X}}_{nl}^{2}=0$, can be described by
${\cal L}\supseteq\int d^{4}\theta\textbf{X}_{nl}^{\dagger}\textbf{X}_{nl}+\left[\int d^{2}\theta f_{0}\textbf{X}_{nl}+h.c.\right]~{}.$ (26)
The nilpotent constraints for the chiral Goldstino superfield can be solved to be
$\textbf{X}_{nl}=\frac{\psi_{X}\psi_{X}}{2F_{X}}+\sqrt{2}\theta\psi_{X}+\theta^{2}F_{X}~{},$ (27)
@@
where all the fields are the functions of $y^{\mu}\equiv x^{\mu}+i\theta\sigma^{\mu}\bar{\theta}$. The lowest component of $\textbf{X}_{nl}$ is eliminated by the nilpotent constraints $\textbf{X}_{nl}^{2}=0$ and is no longer an independent complex scalar field. This leaves only the Goldstino field $\psi_{X}$ and an independent auxiliary field $F_{X}$ in $\textbf{X}_{nl}$ superfield. The Goldstino field $\psi_{X}$, which is the fermionic component of $\textbf{X}_{nl}$, eventually mixes with the gravitino to give it a mass through the super-Higgs mechanism. The lowest component of $\textbf{X}_{nl}$, which is denoted as ${X}_{nl}$ for the Goldstino bilinear ${\psi_{X}\psi_{X}}$, obviously satisfies ${X}_{nl}^{2}=0$ due to its anti-commuting Grassman variable nature. The chiral superfield $\textbf{X}_{nl}$ can be seen as the long-distance expression of the microscopic superfield X that satisfies $\bar{D}^{\dot{\alpha}}{\cal J}_{\alpha\dot{\alpha}}=D_{\alpha}\textbf{X}$ for Ferrara-Zumino multiplet ${\cal J}_{\alpha\dot{\alpha}}$ KS , which can be used even when SUSY is spontaneously broken and can be followed consistently along the renormalization group flow down to the deep IR. As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $\textbf{X}_{nl}$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent Goldstino superfield. From the perspective of low energy effective field theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for discussions on modulus stabilization and vacuum uplifting.
@@
When coupled to supergravity, the most general form (at the two-derivative level) for the action in superspace is given by
@@
$K=\textbf{X}_{nl}^{\dagger}\textbf{X}_{nl}~{},~{}~{}~{}W=f_{0}\textbf{X}_{nl}+c_{1}^{3}~{},$ (29)
where without loss of generality we take $f_{0}$ and $c_{1}$ to be real. Integrating out the auxillary field $F_{X}$ gives $F_{X}=-(W_{\textbf{X}_{nl}})^{*}=-f_{0}$. To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
$K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}~{},$
$W$ $\supseteq$ $\left(f_{0}\textbf{X}_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
The supergravity scalar potential can be derived from
@@
where the Kahler-covariant derivative of the superpotential is
@@
$K^{{\alpha}\bar{\beta}}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}\textbf{X}_{nl}}{3}\\ -\frac{i(2\Im\tau)k_{X}\textbf{X}^{\dagger}_{nl}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}(\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl})}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
Here the indices $\alpha,\beta\sim(\tau,\textbf{X}_{nl})$ and the nilpotency of $\textbf{X}_{nl}$ is used in the calculation.
The scalar potential for the modulus field $\tau$ and $X_{nl}$ can be calculated as
@@
The constraint $X_{nl}^{2}=0$ for the lowest component of $\textbf{X}_{nl}$ is used to eliminate some of the high order $X_{nl}$ terms.
We also use the following expression for the derivative of Dedekind $\eta$ function
@@
The generic complicate form of the scalar potential may obscure the discussions on the stabilization mechanism. Therefore, we would like to concentrate on its typical simplification forms
@@
    When $k_{X}=0$, the scalar potential simplifies significantly.
    The potential can then be written as
@@
    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right\}~{}$
    $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{},$
    where
@@
    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{}.$ (42)
    Note the constraints $X_{nl}^{2}=({X}^{*}_{nl})^{2}=0$ are used in the scalar potential.
    In this scenario, the modulus VEV can be determined by the extreme conditions of the scalar potential
@@
    and the minimum condition can be determined by the positive definiteness of Hessian matrix in terms of
@@
    where $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$. It was argued in target:space ; Ding:2024neh ; Higaki:2024pql that the modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensure that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
@@
    We then determine the extreme conditions
@@
    for $X_{nl}\equiv\rho e^{i\theta}$. We have the following discussions for some non-generic simplified scenarios:
    • Case $\rho=0$:
      This corresponds to $X_{nl}=0$, reducing to the case with only a single $\tau$ modulus field, which has already been extensively discussed target:space ; NPP:2201.02020 .
    • Case $f_{0}=0$:
      The extrema of the scalar potential are located at the parameter points with $X_{nl}=0$. Again, the modulus VEV of the corresponding extrema is the same as the case with only a single $\tau$ modulus field.
    • Case $c_{1}=0$:
      The extrema of the scalar potential are also located at the parameter point with $X_{nl}=0$, and the corresponding potential energy is
      $\left.V\right|_{X_{nl}=0}=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left|{H(\tau)}\right|^{2}f_{0}^{2}\geq 0~{},$ (46)
      for values of $\tau$ within the fundamental region. To obtain a vanishing (very tiny positive) vacuum energy $\Lambda^{4}\sim 10^{-120}M_{Pl}^{4}\approx 0$, we can find the parameter points that are solutions of ${H(\tau)}=0$ and also satisfy the minimum condition (with respect to $\tau$ and $\tau^{*}$) with $\tilde{W}_{M}=0$. It can be checked that solutions of ${H(\tau)}=0$ can also satisfy the minimum condition $\tilde{W}_{M}=0$, which will be clear shortly. Therefore, the stabilized modulus VEV is located at the solutions of ${H(\tau)}=0$ with (almost) vanishing vacuum energy in this case.
    We now focus on the general case with $f_{0},c_{1},X_{nl}\neq 0$. It can be calculated that the extreme condition for $\theta$ is
    $\frac{\partial}{\partial\theta}V=-\frac{2f_{0}c_{1}^{3}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\rho\sin\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]=0.$ (47)
    First, we should note that, as $[(2\Im\tau)|\eta(\tau)|^{4}]^{-1}$ diverges when $\Im\tau=\infty$, the $\tau=i\infty$ point is not a solution of the extreme condition for $\theta$.
    For generic $\theta$ (with $\theta\neq 0,\pi$ to render $X_{nl}$ not real) in the case $f_{0},c_{1},\rho\neq 0$, the extreme condition for $\theta$ reduces to the constraint
    $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{}.$ (48)
    After substituting this constraint in the extreme condition for $\rho$, we have
    $\frac{\partial}{\partial\rho}V=4\rho f_{0}^{2}\frac{\left|H(\tau)\right|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}=0~{}.$ (49)
    Such an extreme condition for $\rho$ can be satisfied with
    $\left|H(\tau)\right|^{2}=0~{}.$ (50)
    Thus, we have the extreme conditions for the case with $f_{0},c_{1}\neq 0$ and a non-vanishing non-real $X_{nl}$
    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{}.$ (51)
    Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes. Therefore, the presence of a non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe in this case.
    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires
    $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0,~{}~{}{\rm or}~{}~{}~{}~{}{\cal P}(j(\tau))=0.$ (52)
    Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively. As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, any complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all unless we specify the concrete form of ${\cal P}(j(\tau))$.
    Then, we would like to focus on the constraint
@@
    which satisfies $H(\tau)=0$. Such a constraint reduces to
@@
    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{}.$
    So $H^{\prime}(\tau)=0$ can trivially hold when $H(\tau)=0$.
@@
      Unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$. Similarly, when $m\neq 0$ and $n=0$, we can deduce that $m\in 2\mathbb{Z}$ should be negative. For $m\neq 0$ and $n\neq 0$, similar deduction can give
      ${\cal P}(j(\tau))=C_{0}^{\prime}\left[j(\tau)\right]^{-n/3}\left[{j(\tau)-1728}\right]^{-m/2}~{},$ (58)
      which also requires both $m\in 2\mathbb{Z}$ and $n\in 3\mathbb{Z}$ being negative to guarantee ${\cal P}(j(\tau))$ being a polynomial of $j(\tau)$. As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
@@
      With the constraints $m=n=0$, the condition (55) reduces to
@@
      As $H(\tau)=0$, this constant should be zero.
    Therefore, the condition (55) requires a vanishing $H(\tau)$. For the condition $j^{\prime}(\tau)=0$, from the derivative of the Klein $j$-function
@@
    we can see that its solutions are $\omega$ and $i$, leading to no new additional constraints. Therefore, the extreme conditions (51) can reduce to the requirement of vanishing $H(\tau)$.
    Before we conclude this scenario, it should be noted that $\theta=0,\pi$, which corresponds to real $X_{nl}=\pm\rho$ with positive $\rho$, are also solutions of the the extreme condition (47) for $\theta$. In this case, the scalar potential reduces to
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (62)
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right.$
    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}\rho^{2}\pm 2f_{0}c_{1}^{3}\rho+\rho^{2}c_{1}^{6}\frac{}{}\right)~{}$
    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right\}~{}$
    $+$ $\frac{\rho^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{}.$
    The scalar potential, which is a quadratic polynomial of $\rho$, can be written as
    $V=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left[C_{1}\rho^{2}\pm C_{2}\rho+C_{3}\right]~{},$ (63)
    where
    $C_{1}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}(f_{0}^{2}+c_{1}^{6})-2\left|H(\tau)\right|^{2}c_{1}^{6}~{},$
    $C_{2}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}2f_{0}c_{1}^{3}-4\left|H(\tau)\right|^{2}f_{0}c_{1}^{3}~{},$
    $C_{3}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}\left(f_{0}^{2}-3c_{1}^{6}\right)~{}.$ (64)
    The scalar potential has a minimum with the corresponding $\rho$ parameter
    $\rho_{0}=\mp\frac{C_{2}}{2C_{1}}~{},$ (65)
    and the potential energy
    $\left.V\right|_{\rho=\rho_{0}}=C_{3}-\frac{C_{2}^{2}}{2C_{1}}~{}.$ (66)
    The modulus VEV is determined by the minimum conditions for the potential (66) with respect to $\tau,\tau^{*}$, and the condition $C_{2}^{2}=2C_{1}C_{3}$ should be established to obtain a vanishing (almost vanishing tiny positive) vacuum energy. It can be checked from previous discussions that solutions of ${H(\tau)}=0$ can also satisfy such minimum conditions.
    Thus, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential are located at the points determined by $H(\tau)=0$, and the corresponding vacuum is of Minkowski type. As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are always the local extrema unless $m=n=0$. Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$. Therefore, other minima, which are determined by (44) to lie at the boundary of the fundamental domain or at the imaginary axis, can also be possible.
    In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices. It can be seen that the potential energies are non-negative when $f_{0}\neq 0$.
    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$. We adopt $\mathcal{P}(j)=1$ for the left panel and $\mathcal{P}(j)=j-4.5764069879884376\times 10^{-11}$ for the right panel.
@@
    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the constraint $X_{nl}^{2}=0$. Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below.
    The scalar potential can be simplified to